# Non-Hermitian Fabry–Pérot Resonances

### v1

@@ Abstract.
We characterise non-Hermitian Fabry–Pérot resonances in high-contrast resonator systems and study the properties of their associated resonant modes from continuous differential models. We consider two non-Hermitian effects: the exceptional point degeneracy and the skin effect induced by imaginary gauge potentials. Using the propagation matrix formalism, we characterise these two non-Hermitian effects beyond the subwavelength regime. This analysis allows us to (i) establish the existence of exceptional points purely from radiation conditions and to (ii) prove that the non-Hermitian skin effect applies uniformly across resonant modes, yielding broadband edge localisation.
@@ 1. Introduction and motivation
In this paper, we consider continuous models for systems of non-Hermitian high-contrast resonators and characterise their Fabry–Pérot resonances. In the subwavelength regime, these scattering resonances (also called resonant frequencies) and effects obtained coincide with those derived earlier through the capacitance matrix formalism [ammari.davies.ea2024Functional]. Two non-Hermitian effects are studied beyond the subwavelength regime: the approximate and exact exceptional points in systems of resonators and the non-Hermitian skin effect induced by imaginary gauge potentials.
@@
Exceptional points typically occur when a system’s resonant frequencies and associated eigenmodes simultaneously coalesce as the system parameters are varied. In terms of finite, discrete systems described by a matrix, they correspond to parameter configurations where the system ceases to be diagonalisable. They are a unique feature of non-Hermitian systems and offer promising applications to enhance sensing due to the higher-order dependence on perturbations of systems at exceptional points [exceptpoint2, hodaei.hassan.ea2017Enhanced, vollmer.arnold.ea2008Single].
@@
The non-Hermitian skin effect is the phenomenon whereby – as a consequence of non-reciprocity – a large proportion of the bulk eigenmodes of a non-Hermitian system are localised at one edge of the system. It allows for unidirectional wave control, offering wide-ranging applications in topological photonics, phononics and other condensed matter systems [hatano.nelson1996Localization, yokomizo, rivero.feng.ea2022Imaginary].
@@
In the first part of the paper, we consider resonator systems with potentially complex wave speeds inside the resonators (in the non-subwavelength regime) and characterise their exceptional points in terms of a mismatch between the order of the zero $\omega_{0}$ of the characteristic determinant obtained from the total propagation matrix and the dimension of the space of solutions of the scattering resonance problem at $\omega_{0}$. By doing so, we recover the results obtained in the subwavelength regime that are based on the non-diagonalisability of the capacitance matrix associated with the resonator system [exceptpoint1, exceptpoint2], which requires the use of complex wave speeds. Both definitions coincide in the subwavelength regime. They also coincide with the fact that a resonant frequency for which exceptional points occur is a pole of the Green function associated to the system that is of order higher than one. This can be seen in the non-subwavelength regime from the fact that the Green function, as the fundamental solution of (4.1), can be expressed in terms of propagation matrices [SANCHEZSOTO2012191]. In the subwavelength regime, the (discrete) Green function at a frequency $\omega$ is nothing more than the inverse of $\omega^{2}I-\tilde{\mathcal{C}}(0)$ with $\tilde{\mathcal{C}}(0)$ being the capacitance matrix of the resonator system [ammari.davies.ea2024Functional]. Most of the exceptional points discussed in this part are approximate in the sense that they are achieved by ensuring the coalescence of the leading-order behaviour of resonances as the contrast material parameter goes to zero. Nevertheless, by changing the radiation conditions into parity-time symmetric ones, we show analytically that such systems with balanced energy gain and loss have exact exceptional points and demonstrate numerically the square-root coalescence to parameters around both the approximate and exact exceptional points.
@@
The second part of this paper is devoted to the study of the non-Hermitian skin effect in the non-subwavelength regime. Using again propagation matrices for the system and in this non-reciprocal case their symmetrisation, we characterise the scattering resonances of the Helmholtz equation with imaginary gauge potential inside the resonators and show that the skin effect holds even for non-subwavelength modes. The eigenmodes of the resonator system still exhibit a uniform decay that is independent of frequency. Nevertheless, compared to the subwavelength regime, they are not constant inside the resonators; instead, they may highly oscillate. Note that in [ammari.barandun.ea2023NonHermitian] the occurrence of the non-Hermitian skin effect in the presence of an imaginary gauge potential is proved using Toeplitz matrix theory for a discrete approximation of the continuous model (which is based on differential equations) in terms of the gauge capacitance matrix that holds only in the subwavelength regime. In this paper we prove the occurrence of the non-Hermitian skin effect for the continuous problem both in the subwavelength and non-subwavelength regimes.
@@
In fact, our results in this paper constitute the first results on non-Hermitian scattering resonances for systems of resonators beyond the subwavelength regime. They give an analytical and numerical framework for computing the scattering resonances in two non-Hermitian settings. In particular, they show that non-Hermitian phenomena (exceptional degeneracies and accumulation of eigenmodes at one edge of the system) that are known to occur in the subwavelength regime hold also in the non-subwavelength regime. The use of the new discrete approximation in terms of the generalised capacitance matrix in the non-subwavelength regime that is introduced in [pm2] allows for a precise description of the resonances in the non-Hermitian settings considered here, how many there are, and the asymptotic properties of their corresponding resonant states. It is expected that such a characterisation holds in three dimensions, which would also lead to the study of non-Hermitian phenomena in three-dimensional systems beyond the subwavelength regime directly from the continuous model.
@@
Our paper is organised as follows. In Section˜2, we introduce the scattering resonance problem and its capacitance matrix formulation that hold beyond the subwavelength regime. Then, in Section˜3, we elucidate the variety of exceptional points (exact and asymptotic) that arise in the non-Hermitian setting. Namely, in Section˜3.3 we characterise exceptional points induced by leading-order Hermiticity in the form of complex material parameters. In Section˜3.4, we demonstrate how exceptional points may be obtained from the non-Hermiticity due to outgoing radiation. Moreover, in Section˜3.5, we demonstrate that a $\mathcal{PT}$-symmetric modified radiation condition allows for explicit exceptional points. Section˜4 is devoted to the study of the scattering resonances in non-reciprocal systems and to the proof of the non-Hermitian skin effect in such systems by means of a symmetrisation approach. By combining our approaches in this paper with those in [junshan1, ammari.barandun.ea2025Generalized, barandun2023], we expect to generalise our results in this paper to periodic infinite (for not only the exceptional point degeneracy but also the Dirac degeneracy) and semi-infinite (for the non-Hermitian skin effect) structures.
@@ 2. Setting and capacitance matrix formalism
the set of resonators.
@@
As $\delta\to 0$, the characteristic determinant converges uniformly on any compact set in $\omega$. Moreover, the following characterisation of the zeros of $f(\omega;0)$ holds; see [pm1, Lemma 3.5].
@@
This result justifies the transmission comb observed in Figure˜1 for a dimer system ($N=2$) with $\ell_{1}=\ell_{2}=s=1$ and $v=v_{i}=1$. The threefold zeros of $f(\omega;0)$ around $\omega=0,\pi$ split into three distinct zeros with small imaginary parts for small $\delta>0$. We let the transmission coefficient $T(\omega)$ for real $\omega$ be defined by
@@
Figure˜1 exhibits the combs with 2N12N-1 peaks for $T(\omega)$ that are typical for Fabry-Pérot-type resonances.
@@
In the asymptotic regime, when $\delta\rightarrow 0$, the leading order behaviour of these scattering resonances is described by discrete capacitance matrix formulations [pm1, pm2]. We let the vectors $\mathfrak{\bm{t}}=(\mathfrak{t}_{1},\cdots,\mathfrak{t}_{2N-1})^{\top}$ and $\bm{t}=(t_{1},\dots,t_{2N-1})^{\top}$ be defined by
@@
The following result is from [pm2].
@@
Note that in the subwavelength regime, when we search for resonant frequencies that are in a neighborhood of zero (i.e., for $\omega_{0}=0$), the generalised capacitance matrix formalism reduces to the study of $\tilde{\mathcal{C}}(0)$, which was first introduced in [feppon.cheng.ea2023Subwavelength]. It is also worth noting that if the wave speeds $v_{i}$ are all real positive, then $\tilde{\mathcal{C}}(0)$ is always diagonalisable, whereas this is not necessarily the case for general $\omega_{0}\in\{\omega\mid f(\omega;0)=0\}$. The diagonalisability of $\tilde{\mathcal{C}}(0)$ when the wave speeds are all positive follows from the fact that $V\tilde{\mathcal{C}}(0)$, where $V$ is the diagonal positive definite matrix given by $V:=\mathrm{diag}(t_{1},t_{3},\dots,t_{2N-1})$, is a real symmetric matrix and hence $\tilde{\mathcal{C}}(0)$ is similar to $\sqrt{V^{-1}}V\tilde{\mathcal{C}}(0)\sqrt{V^{-1}}$.
@@ 3.1. Exceptional points and Green’s function
The key property of exceptional points is that degenerate resonances exhibit higher-order sensitivity with respect to generic perturbations [kato1995Perturbation, Chapter II, 1.2]. In particular, for an exceptional point of order two, the two degenerate resonant frequencies $\omega_{\pm}$ perturb as
@@
where the Green function is given by (cf. [kato1995Perturbation, Sec. 2.3, Chap. 3] and the references therein)
@@ 3.2. Exceptional points in the limit $\delta=0$
This behavior is illustrated in Figure˜2 for a perturbed dimer system with resonator lengths $\ell_{1}=1$ and $\ell_{2}=1.1$, focusing on resonances near 0 and $\pi$. Around 0, all eigenmodes interact, resulting in $\sqrt{\delta}$-order convergence. However, the situation is more nuanced around $\pi$: due to $s=\ell_{1}=1$, the resonant mode of the first resonator interacts with that of the gap between the resonators, again yielding $\sqrt{\delta}$-order convergence to the limiting mode. In contrast, the resonant mode of the second resonator does not interact with the eigenmode of the gap between the resonators due to the length mismatch, resulting in $\delta$-order convergence.
@@ 3.3. Approximate exceptional points from leading-order non-Hermiticity
In this subsection, we examine the exceptional points arising from leading-order non-Hermiticity, induced by complex material parameters, similar to the analysis in [exceptpoint1].
@@
Here, $\omega_{3},\omega_{2}$ denote the two subwavelength resonant frequencies with positive real parts as $\delta\to 0$. In fact, this coalescence dichotomy is observed in Figure˜3(a) for the trimer system at and away from the exceptional point after tuning $\theta$ (in particular, we compare the coalescence behaviour for $\theta=\pi/4$ and $\theta=0$).
For small $\delta>0$, the higher order coalescence of the resonant frequencies acts as an approximate exceptional point. In particular, the eigenvalues of $\tilde{\mathcal{C}}(0)$ and thus the leading-order behaviour of $\omega_{1},\omega_{2}$ display higher-order sensitivity to system parameters. This is demonstrated in Figure˜4(a) where, as the first resonator size $\ell_{1}$ is varied around $1$, the resonant frequencies $\omega_{1},\omega_{2}$ demonstrate the characteristic higher-order sensitivity to perturbations. At this point, we would also like to note that, while this constitutes only an approximate exceptional point, the discussion from the following subsection would suggest that there is an exact exceptional point in a system-parameter neighbourhood.
@@ 3.4. Approximate exceptional points from radiation losses
determining the critical values $\theta_{c}=3\pm 2\sqrt{2}$. At these critical values, the ansatz $\omega_{\pm}=\pi+\lambda_{\pm}\delta+\mathcal{O}(\delta^{\frac{3}{2}})$ suggests a coalescence of $\omega_{\pm}$ with order $\mathcal{O}(\delta^{3/2})$, yielding an approximate exceptional point. This coalescence behaviour is demonstrated in Figure˜3. In Figure˜4(b), we observe that the approximate exceptional point indeed also displays the expected square-root sensitivity to perturbations in $\theta$ around the critical values.

### v2

@@ Abstract.
We characterise non-Hermitian Fabry–Pérot resonances in high-contrast resonator systems and study the properties of their associated resonant modes from continuous differential models. We consider two non-Hermitian effects: the exceptional point degeneracy and the skin effect induced by imaginary gauge potentials. Using a propagation matrix formalism, we characterise these two non-Hermitian effects beyond the subwavelength regime. This analysis allows us to (i) establish the existence of exceptional points purely from radiation conditions and to (ii) prove that the non-Hermitian skin effect applies uniformly across resonant modes, yielding broadband edge localisation.
@@ 1. Introduction and motivation
In this paper, we consider continuous models for systems of non-Hermitian high-contrast resonators and characterise their Fabry–Pérot resonances. In the subwavelength regime, these scattering resonances (also called resonant frequencies) and effects obtained coincide with those derived earlier through the capacitance matrix formalism [undefe]. Two non-Hermitian effects are studied beyond the subwavelength regime: the approximate and exact exceptional points in systems of resonators and the non-Hermitian skin effect induced by imaginary gauge potentials.
@@
Exceptional points typically occur when a system’s resonant frequencies and associated eigenmodes simultaneously coalesce as the system parameters are varied. In terms of finite, discrete systems described by a matrix, they correspond to parameter configurations where the system ceases to be diagonalisable. They are a unique feature of non-Hermitian systems and offer promising applications to enhance sensing due to the higher-order dependence on perturbations of systems at exceptional points [undefg, undefn, undefu].
@@
The non-Hermitian skin effect is the phenomenon whereby – as a consequence of non-reciprocity – a large proportion of the bulk eigenmodes of a non-Hermitian system are localised at one edge of the system. It allows for unidirectional wave control, offering wide-ranging applications in topological photonics, phononics and other condensed matter systems [undefm, undefv, undefr].
@@
In the first part of the paper, we consider resonator systems with potentially complex wave speeds inside the resonators (in the non-subwavelength regime) and characterise their exceptional points in terms of a mismatch between the order of the zero $\omega_{0}$ of the characteristic determinant obtained from the total propagation matrix and the dimension of the space of solutions of the scattering resonance problem at $\omega_{0}$. By doing so, we recover the results obtained in the subwavelength regime that are based on the non-diagonalisability of the capacitance matrix associated with the resonator system [undeff, undefg], which requires the use of complex wave speeds. Both definitions coincide in the subwavelength regime. They also coincide with the fact that a resonant frequency for which exceptional points occur is a pole of the Green function associated to the system that is of order higher than one. This can be seen in the non-subwavelength regime from the fact that the Green function, as the fundamental solution of (4.1), can be expressed in terms of propagation matrices [undefs]. In the subwavelength regime, the (discrete) Green function at a frequency $\omega$ is nothing more than the inverse of $\omega^{2}I-\tilde{\mathcal{C}}(0)$ with $\tilde{\mathcal{C}}(0)$ being the capacitance matrix of the resonator system [undefe]. Most of the exceptional points discussed in this part are approximate in the sense that they are achieved by ensuring the coalescence of the leading-order behaviour of resonances as the contrast material parameter goes to zero. Nevertheless, by changing the radiation conditions into parity-time symmetric ones, we show analytically that such systems with balanced energy gain and loss have exact exceptional points and demonstrate numerically the square-root coalescence to parameters around both the approximate and exact exceptional points.
@@
The second part of this paper is devoted to the study of the non-Hermitian skin effect in the non-subwavelength regime. Using again propagation matrices for the system and in this non-reciprocal case their symmetrisation, we characterise the scattering resonances of the Helmholtz equation with imaginary gauge potential inside the resonators and show that the skin effect holds even for non-subwavelength modes. The eigenmodes of the resonator system still exhibit a uniform decay that is independent of frequency. Nevertheless, compared to the subwavelength regime, they are not constant inside the resonators; instead, they may highly oscillate. Note that in [undefa] the occurrence of the non-Hermitian skin effect in the presence of an imaginary gauge potential is proved using Toeplitz matrix theory for a discrete approximation of the continuous model (which is based on differential equations) in terms of the gauge capacitance matrix that holds only in the subwavelength regime. In this paper we prove the occurrence of the non-Hermitian skin effect for the continuous problem both in the subwavelength and non-subwavelength regimes.
@@
In fact, our results in this paper constitute the first results on non-Hermitian scattering resonances for systems of resonators beyond the subwavelength regime. They give an analytical and numerical framework for computing the scattering resonances in two non-Hermitian settings. In particular, they show that non-Hermitian phenomena (exceptional degeneracies and accumulation of eigenmodes at one edge of the system) that are known to occur in the subwavelength regime hold also in the non-subwavelength regime. The use of the new discrete approximation in terms of the generalised capacitance matrix in the non-subwavelength regime that is introduced in [undefh] allows for a precise description of the resonances in the non-Hermitian settings considered here, how many there are, and the asymptotic properties of their corresponding resonant states. It is expected that such a characterisation holds in three dimensions, which would also lead to the study of non-Hermitian phenomena in three-dimensional systems beyond the subwavelength regime directly from the continuous model.
Moreover, while this paper considered simple finite or infinite periodic geometries, another open question is the characterisation of bandgaps for spatially disordered resonator arrays beyond the subwavelength regime. We expect that adapting [undefi] to the propagation matrix framework used in this work should allow for the guaranteed existence of bandgaps from local properties.
@@
Our paper is organised as follows. In Section˜2, we introduce the scattering resonance problem and its capacitance matrix formulation that hold beyond the subwavelength regime. Then, in Section˜3, we elucidate the variety of exceptional points (exact and asymptotic) that arise in the non-Hermitian setting. Namely, in Section˜3.3 we characterise exceptional points induced by leading-order Hermiticity in the form of complex material parameters. In Section˜3.4, we demonstrate how exceptional points may be obtained from the non-Hermiticity due to outgoing radiation. Moreover, in Section˜3.5, we demonstrate that a $\mathcal{PT}$-symmetric modified radiation condition allows for explicit exceptional points. Section˜4 is devoted to the study of the scattering resonances in non-reciprocal systems and to the proof of the non-Hermitian skin effect in such systems by means of a symmetrisation approach. By combining our approaches in this paper with those in [undefq, undefc, undefb], we expect to generalise our results in this paper to periodic infinite (for not only the exceptional point degeneracy but also the Dirac degeneracy) and semi-infinite (for the non-Hermitian skin effect) structures.
@@ 2. Setting and capacitance matrix formalism
the set of resonators; see Figure˜1.
$\ell_{1}$$x_{1}^{\operatorname{{\mathsf{L}}}}$$x_{1}^{\operatorname{{\mathsf{R}}}}$$\ell_{2}$$x_{2}^{\operatorname{{\mathsf{L}}}}$$x_{2}^{\operatorname{{\mathsf{R}}}}$$\ell_{3}$$x_{3}^{\operatorname{{\mathsf{L}}}}$$x_{3}^{\operatorname{{\mathsf{R}}}}$$\ell_{4}$$x_{4}^{\operatorname{{\mathsf{L}}}}$$x_{4}^{\operatorname{{\mathsf{R}}}}$$s_{1}$$s_{2}$$s_{3}$$x_{N-1}^{\operatorname{{\mathsf{L}}}}$$x_{N-1}^{\operatorname{{\mathsf{R}}}}$$\ell_{N-1}$$x_{N}^{\operatorname{{\mathsf{L}}}}$$x_{N}^{\operatorname{{\mathsf{R}}}}$$\ell_{N}$$s_{N-1}\operatorname{\mathbb{R}}$
Figure 1. A one-dimensional chain of $N$ disjoint resonators.
@@
As $\delta\to 0$, the characteristic determinant converges uniformly on any compact set in $\omega$. Moreover, the following characterisation of the zeros of $f(\omega;0)$ holds; see [undefo, Lemma 3.5].
@@
This result justifies the transmission comb observed in Figure˜2 for a dimer system ($N=2$) with $\ell_{1}=\ell_{2}=s=1$ and $v=v_{i}=1$. The threefold zeros of $f(\omega;0)$ around $\omega=0,\pi$ split into three distinct zeros with small imaginary parts for small $\delta>0$. We let the transmission coefficient $T(\omega)$ for real $\omega$ be defined by
@@
Figure˜2 exhibits the combs with 2N12N-1 peaks for $T(\omega)$ that are typical for Fabry-Pérot-type resonances.
@@
In the asymptotic regime, when $\delta\rightarrow 0$, the leading order behaviour of these scattering resonances is described by discrete capacitance matrix formulations [undefo, undefh]. We let the vectors $\mathfrak{\bm{t}}=(\mathfrak{t}_{1},\cdots,\mathfrak{t}_{2N-1})^{\top}$ and $\bm{t}=(t_{1},\dots,t_{2N-1})^{\top}$ be defined by
@@
The following result is from [undefh].
@@
Note that in the subwavelength regime, when we search for resonant frequencies that are in a neighborhood of zero (i.e., for $\omega_{0}=0$), the generalised capacitance matrix formalism reduces to the study of $\tilde{\mathcal{C}}(0)$, which was first introduced in [undefk]. It is also worth noting that if the wave speeds $v_{i}$ are all real positive, then $\tilde{\mathcal{C}}(0)$ is always diagonalisable, whereas this is not necessarily the case for general $\omega_{0}\in\{\omega\mid f(\omega;0)=0\}$. The diagonalisability of $\tilde{\mathcal{C}}(0)$ when the wave speeds are all positive follows from the fact that $V\tilde{\mathcal{C}}(0)$, where $V$ is the diagonal positive definite matrix given by $V:=\mathrm{diag}(t_{1},t_{3},\dots,t_{2N-1})$, is a real symmetric matrix and hence $\tilde{\mathcal{C}}(0)$ is similar to $\sqrt{V^{-1}}V\tilde{\mathcal{C}}(0)\sqrt{V^{-1}}$.
@@ 3.1. Exceptional points and Green’s function
The key property of exceptional points is that degenerate resonances exhibit higher-order sensitivity with respect to generic perturbations [undefp, Chapter II, 1.2]. In particular, for an exceptional point of order two, the two degenerate resonant frequencies $\omega_{\pm}$ perturb as
@@
where the Green function is given by (cf. [undefp, Sec. 2.3, Chap. 3] and the references therein)
@@ 3.2. Exceptional points in the limit $\delta=0$
This behavior is illustrated in Figure˜3 for a perturbed dimer system with resonator lengths $\ell_{1}=1$ and $\ell_{2}=1.1$, focusing on resonances near 0 and $\pi$. Around 0, all eigenmodes interact, resulting in $\sqrt{\delta}$-order convergence. However, the situation is more nuanced around $\pi$: due to $s=\ell_{1}=1$, the resonant mode of the first resonator interacts with that of the gap between the resonators, again yielding $\sqrt{\delta}$-order convergence to the limiting mode. In contrast, the resonant mode of the second resonator does not interact with the eigenmode of the gap between the resonators due to the length mismatch, resulting in $\delta$-order convergence.
@@ 3.3. Approximate exceptional points from leading-order non-Hermiticity
In this subsection, we examine the exceptional points arising from leading-order non-Hermiticity, induced by complex material parameters, similar to the analysis in [undeff].
@@
Here, $\omega_{3},\omega_{2}$ denote the two subwavelength resonant frequencies with positive real parts as $\delta\to 0$. In fact, this coalescence dichotomy is observed in Figure˜4(a) for the trimer system at and away from the exceptional point after tuning $\theta$ (in particular, we compare the coalescence behaviour for $\theta=\pi/4$ and $\theta=0$).
For small $\delta>0$, the higher order coalescence of the resonant frequencies acts as an approximate exceptional point. In particular, the eigenvalues of $\tilde{\mathcal{C}}(0)$ and thus the leading-order behaviour of $\omega_{1},\omega_{2}$ display higher-order sensitivity to system parameters. This is demonstrated in Figure˜5(a) where, as the first resonator size $\ell_{1}$ is varied around $1$, the resonant frequencies $\omega_{1},\omega_{2}$ demonstrate the characteristic higher-order sensitivity to perturbations. At this point, we would also like to note that, while this constitutes only an approximate exceptional point, the discussion from the following subsection would suggest that there is an exact exceptional point in a system-parameter neighbourhood.
@@ 3.4. Approximate exceptional points from radiation losses
determining the critical values $\theta_{c}=3\pm 2\sqrt{2}$. At these critical values, the ansatz $\omega_{\pm}=\pi+\lambda_{\pm}\delta+\mathcal{O}(\delta^{\frac{3}{2}})$ suggests a coalescence of $\omega_{\pm}$ with order $\mathcal{O}(\delta^{3/2})$, yielding an approximate exceptional point. This coalescence behaviour is demonstrated in Figure˜4. In Figure˜5(b), we observe that the approximate exceptional point indeed also displays the expected square-root sensitivity to perturbations in $\theta$ around the critical values.